Division of Spill Prevention and Response

Breadcrumbs

Chronology

April

March

February


April 10, 2012 (final update)

Spill responders from Alaska Clean Seas (ACS), Penco and CCI used heavy equipment, such as loaders, excavators, trimmers, skid-steers, hydraulic hammers and snow machines with trailers, to remove the spill material from the heaviest contaminated zones. Hand tools were used to recover drilling mud from the areas of lightest contamination. Cleanup crews removed approximately 6,286 cubic yards of spilled material from the snow-covered tundra. In addition, 2,402 cubic yards of down-hole material and 116,928 gallons of water-based mud and freshwater from the steam units were recovered from the drilling pad during the initial response.

Repsol hired a surveying contractor to assist in calculating the off-pad impacted area and the volume of the material released to the tundra. The contractor estimated that 21,114 bbls of down-hole material were released to the tundra during the initial blow-out. In addition, they calculated the total impacted area as 23.75 acres. An area of 16.76 acres of lightly-misted material was determined to be unrecoverable, and no cleanup occurred in this zone.

On April 1, 2012, a group from the Kuukpikmiut Subsistence Oversight Panel, Inc. (KSOP) and from the village of Nuiqsut visited the site to see how the cleanup had progressed. The group was satisfied with the work that had been done and said that they did not believe that any further cleanup needed to be performed.

On April 5, 2012, the final site inspection team, composed of representatives from the North Slope Borough, KSOP, ADEC, ADNR and Repsol, made the final site inspection. The team agreed that no further cleanup actions were required.

ADEC's field monitors were demobilized from the site on April 6, 2012.

 


March 18, 2012

The well is plugged and controlled. The Unified Command has shifted the response management into the project phase and has demobilized the ICP in Anchorage.

The cleanup plan has been approved and is flexible so that information gained during the final delineation operations will be used to drive the cleanup actions.

The off-pad clean-up area has been divided into four zones. Delineation of two of the four zones is complete. Delineation of the final two zones will be completed today when the weather permits. Due to the snow cover over the area it is difficult to identify the edges of the plume. When temperatures allow, a skid steer with broom and snow machines will be used to clear away the layer of clean snow which has accumulated over the top of the deposition.

Cold temperatures continue to slow the work as field crews are on a weather hold until the temperature moderates to -45°F, when they can perform hand work, and -35°F, when machines can be operated.
The temperature on site this morning was -41°F with a 10 mph wind.


March 16, 2012

Two cement plugs have been set in the well and pressure tested. Although the well plug and abandonment efforts are ongoing, it has been determined that it is safe to begin the cleanup activities. Field crews are ready to begin the site delineation, which will give information critical for decision making on appropriate removal actions. The information gained during the delineation operation will be used to finalize the cleanup plan. The temperature on site this morning was -50°F. Field crews are on a weather hold until the temperature moderates to -45°F at which point they will begin the delineation operation.

Repsol reports that a total of 2,189 barrels (91,938 gallons) of liquids (thawed drilling mud and water) have been manifested and shipped offsite to date. Much of the liquid collected is condensed water from the steam used to thaw and clean the rig. A total to 2,363 cubic yards of solid waste (frozen drilling mud and downhole materials) have been manifested and shipped offsite to date.


March 13, 2012

Since the previous Sitrep (20), a coiled tubing unit cleared out a blockage inside the drill pipe. With the blockage removed, a wireline diagnostic well logging was made. The well logging indicated the presence of two obstructions, called bridges, in the well annulus (the space between the drill pipe and the formation). Last night, the well control crew attempted to move the drill pipe to clear the bridges and establish circulation between the inside of the drill pipe and the annulus. The attempt failed and they were unable to circulate fluid through the well.

Repsol has made the decision to plug and abandon (P&A) the well, thereby achieving well control. Repsol is finalizing the well control plan for review and approval by the Alaska Oil and Gas Conservation Commission (AOGCC). Upon approval of the well control plan by AOGCC, Repsol will begin the well control operation.

Repsol reports that a total of 2,093 barrels (87,906 gallons) of liquids have been manifested and shipped offsite. Liquids collected may include drilling mud, water and solids from the well, and condensed water from the steam used to thaw and clean the rig. A total of 2,299 cubic yards of solid waste (frozen drilling mud and downhole materials) have been manifested and shipped offsite.


March 06, 2012

The primary actions being taken now continue to be work towards successful well control.

A coiled tubing unit and crane are on site and is being rigged to clean out the blockage. The first step is to pump a heavy mud into the well annulus (the space between the drill pipe and the formation) to stabilize it. With the annulus stabilized, Repsol will run the coil tube down the drill stem and attempt to melt and wash out the obstruction. Once the coil tube has cleared the obstruction and the coiled tubing crew get down towards the foot near the drill bit, they will start pumping in a heavy completion fluid. If the coiled tubing effort is successful and the blockage is cleared, Repsol will evaluate the condition of the well and work to finalize the well-kill plan.

Since the previous Sitrep (19), Repsol reports that 58 barrels (2,436 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal, bringing the total to 1,858 barrels (78,036 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid collected is condensed water from the steam used to thaw and clean the rig. Another 569 cubic yards of solid waste (frozen drilling mud) were manifested and shipped offsite, bringing the total to 1,411 cubic yards of solid waste that have been manifested and shipped offsite.


March 05, 2012

No update.


March 04, 2012

No update.


March 03, 2012

Temperatures remain extremely cold. With an ambient temperature of -44°F and wind chill of -68°F, use of equipment outside the drilling rig continues to be limited.

The first well log, made last night, showed a blockage of the drill pipe. A coiled tubing unit and crane have been contracted to clean out the drill pipe, but the temperature must be above -35°F for them to operate. The weather is not forecasted to warm above -35°F until this Tuesday, March 6. The contracted coiled tubing unit is currently working on another job, which will take about one day to complete once the temperatures warm up above -35°F.

Once the coiled tubing unit completes its present job, it will be moved to the Q2 pad, rigged and used to clean out the blockage. Once the blockage is cleared, the wireline crew will run another well logging. If no further problems are discovered, the information gained in the second well logging will be used to finalize the well-kill plan.

Repsol reports that 68 barrels (2,856 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,800 barrels (75,600 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid collected is condensed water from the steam used to thaw and clean the rig. Another 178 cubic yards of solid waste (frozen drilling mud) were manifested and shipped offsite yesterday, bringing the total to 842 cubic yards of solid waste that have been manifested and shipped offsite.

The Unified Command decided to temporarily close the Anchorage Incident Command Post (ICP) until Tuesday morning, March 6, 2012, due to the temperature hold on the wellbore clean out operation. ADEC responders at Palm Pad will remain in place, and ADEC responders working in Anchorage will work out of the local ADEC office until the ICP reopens. If the temperature moderates earlier than currently predicted, the ICP will reopen earlier as well.


March 02, 2012

Response crews continue 24-hour operations to clean and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

The Unified Command (UC) traveled to Nuiqsut for a public meeting yesterday, March 2, 2012. About 30 people attended the meeting at which the UC gave information and answered questions about what had occurred, and current and planned operations.

Temperatures remain extremely cold with a -63°F wind chill, so outside equipment use continues to be limited. However, work inside the rig continues.

Crews have completed the cleaning and drying of the rig's power components and the rig is operational.

A wireline is used to lower instruments down the borehole of the well to gather information on the current condition of the well; a process called well logging. The wireline crew has initiated this process. The information gathered in the well logging will be used for finalization of the well-kill plan.

Repsol reports that 8 barrels (336 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,732 barrels (72,744 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid collected is condensed water from the steam used to thaw and clean the rig. Another 114 cubic-yards of solid waste (frozen drilling mud) were manifested and shipped offsite yesterday bringing the total to 664 cubic-yards of solid waste that have been manifested and shipped offsite.


March 01, 2012

Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

Temperatures remain extremely cold with a -58°F wind chill, so outside equipment continues to be limited. However, work inside the rig continues.

Crews are continuing the cleaning and drying of the rig's power components; critical steps in the eventual re-start of the rig.

The re-start of the rig is one in a series of steps in the progression needed for completion of the well-kill operation. However, due to the need for a very methodical and careful cleaning and drying of the rig's power components, the time of the re-start cannot be determined at this time.

Work on setting up the wireline equipment for well inspection has begun. The wireline is used to lower instruments down the borehole of the well to gather information on the current condition of the well. The information gathered will be used for finalization of the well-kill plan.

Repsol reports that 58 barrels (2,436 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,724 barrels (72,408 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid being collected is condensed water from the steam used to thaw the frozen drilling mud. Another 75 cubic yards of solid waste (frozen drilling mud) were also manifested and shipped offsite yesterday bringing the total to 550 cubic yards of solid waste that have been manifested and shipped offsite.


February 29, 2012

Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

Temperatures remain extremely cold with a -71°F wind chill, so outside equipment continues to be shut down. However, work inside the rig continues.

Lights and electrical heat are energized in parts of the rig. Crews are continuing the cleaning and drying of the rig's power components, a critical step in the eventual re-start of the rig.

The re-start of the rig is one in a series of steps in the progression needed for completion of the well-kill operation. However, due to the need for a very methodical and careful cleaning and drying of the rig's power components, the time of the re-start cannot be determined at this time.

Repsol reports that 232 barrels (9,744 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,666 barrels (69,972 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid being collected is condensed water from the steam used to thaw the frozen drilling mud. Due to the extremely cold temperatures, no solid wastes were manifested or shipped offsite yesterday.


February 28, 2012

Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

The temperature today is -49°F with a 9 mph wind giving a -74°F wind chill, so outside equipment use has been shut down. However, work inside the rig continues.

Both the cold start generator and the #3 generator are up and running. Lights and electrical heat are energized in parts of the rig. Crews are continuing the cleaning and drying of the rig's power components, a critical step in the eventual re-start of the rig.

The re-start of the rig is one in a series of steps in the progression needed for completion of the well-kill operation. However, due to the need for a very methodical and careful cleaning and drying of the rig's power components, the time of the re-start cannot be determined at this time.

Repsol reports that 174 barrels (7,308 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,434 barrels (60,228 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid being collected is condensed water from the steam used to thaw the frozen drilling mud. Another 150 cubic-yards of solid waste were also manifested and shipped offsite yesterday.


February 27, 2012

Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.

Technicians will inspect the cold start generator today. The cold start generator will be the first generator to be started and supplies the power needed for the eventual re-start of the rig. The re-start of the rig is one in a series of steps in the progression needed for completion of the well-kill operation. However, due to the need for a very methodical and careful cleaning and drying of the rig's power components, the time of the re-start cannot be determined at this time.

Repsol reports that 232 barrels (9,744 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,260 barrels (52,920 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid being collected is condensed water from the steam used to thaw the frozen drilling mud. Another one hundred and twenty five cubic-yards of solid waste were also manifested and shipped offsite yesterday.


February 26, 2012

Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.
Repsol reports that another 174 barrels (7,308 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal in the past day, bringing the total to 1,028 barrels (43,176 gallons) of liquids that have been manifested and shipped offsite. Much of the liquid being collected is condensed water from the steam used to thaw the frozen drilling mud. One hundred and twenty five cubic-yards of solid waste were also manifested and shipped offsite yesterday.


February 25, 2012

Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation.
Repsol reports that 116 barrels (4,872 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal yesterday, bringing the total to 854 barrels (35,868 gallons) of liquids that have been manifested and shipped offsite. Fifty cubic-yards of solid waste was also manifested and shipped offsite yesterday.
A public meeting was held in Nuiqsut from 7:00 PM to 11:30 PM last night to address any questions or concerns about the incident. Representatives from ADEC, Repsol, and North Slope Borough were at Nuiqsut for the meeting. The North Slope Borough is holding a workshop this afternoon followed by a public meeting for the community tonight.


February 24, 2012

Response crews continue 24-hour operations to steam thaw and remove drilling mud from the drilling rig in order to access the essential drilling rig components needed for the well-kill operation. Repsol reports that 289 barrels (12,138 gallons) of liquids (thawed drilling mud and water) were shipped offsite for disposal yesterday, bringing the total to 738 barrels (30,996 gallons) of liquids that have been manifested and shipped offsite. One 25-cubic-yard load of solid waste was also manifested and shipped offsite yesterday. A public meeting is scheduled in Nuiqsut from 7:00 PM to 9:00 PM tonight to address any questions or concerns about the incident. Representatives from ADEC, Repsol, and North Slope Borough are at Nuiqsut for the meeting.


February 23, 2012

Response crews continue twenty]four hour operations. Two large heaters pump hot air into the drilling rig for thawing necessary equipment. Crews continue steam thawing and removing drilling mud. Removal remains focused on the waste mud material that blocks access to essential drilling rig components needed for well]kill operations. Complete drilling rig cleanup will occur after the well has been successfully killed and secured. An inspection camera was received onsite and used to evaluate the bore hole integrity, which appears to be in good condition. This will allow the pumping of drilling mud down the well for the final step in the well]kill process. Meanwhile, the hydraulic system was function]tested and no leaks were observed. Repsol reports that 671 bbls of liquids (thawed drilling mud and water) have been manifested and shipped offsite for disposal. Approximately 340 cubic yards of frozen drilling mud waste material is stockpiled on a portion of the ice pad and awaiting transport by trucks for disposal at the designated grind and inject facility.


February 22, 2012

Response crews continue twenty‐four hour operations. Two large heaters pump hot air into the drilling rig for thawing necessary equipment. Steam thawing and removal of drilling mud is focused on spilled material that blocks access to drilling rig components needed for the well‐kill operations. Complete rig cleanup will occur after the well has been successfully killed and secured.
Four supersucker trucks are being used to remove thawed materials. A total of 117 bbls of thawed drilling mud and water have been removed from the site for offsite disposal.
Crews are inspecting drill pipe and effecting repairs where needed. An inspection camera was mobilized to the site to facilitate additional drill pipe inspection. Repsol reported that 200 gallons of hydraulic fluid was lost out of the hydraulic system on the drill rig. The fluid drained to the sump on the rig and will be removed with the drilling mud. Crews continue to make repairs to the hydraulic system.
Construction of the 200‐foot‐by‐200‐foot ice pad extension is complete, and two ten‐man berthing camps will be relocated to this temporary ice pad.


February 21, 2012

Well control response work continues at Q2 pad. Two boilers are operational and providing steam for de‐icing. Portable air heaters are on‐site for general heating of the drilling rig.
Thawing operations remain focused on equipment in the drill cellar that is essential to operating the well. Inspections continue on the hydraulic system as workers de‐ice more of the hydraulic equipment. The gas diverter pipe and one hydraulic hose to the gas diverter valve were found to be damaged. Crews have temporary patches and clamps on‐site but are waiting to repair the damage until more of the critical components are exposed by further de‐icing. Crews installed a pump manifold on the drill rig floor in further preparation for pumping fluid down‐hole into the drill pipe.
A 300‐barrel vacuum truck is on‐site recovering thawed fluids. A supersucker has been ordered to assist in the cleanout of the cellar area. Drilling mud, water, and other material ejected from the well will be trucked to Prudhoe Bay for disposal in an approved injection well.
ADNR has permitted a 200‐foot‐by‐200‐foot temporary extension to the northeast corner of the Q2 pad. This extension will be constructed of ice and will be used to stage additional equipment, such as a warm‐up shack, two small man‐camps, a clinic, etc.


February 20, 2012

Nabors Drilling and Cruz Construction Inc. workers have removed approximately 200 cubic yards of drilling mud and contaminated snow from areas adjacent to the east, north, and west sides of the drilling rig. This will improve access for crews and equipment working to cleanup and repair the rig. Due to the potential danger posed by the gas diverter in the event the well resumes flow, no cleanup will take place on or off the pad to the south of the rig before well control work is complete. Well control workers have set up boilers outside the rig and begun using steam to thaw frozen drilling mud and water from around the drill cellar access doors. Thawing operations will first focus on gaining access to hydraulic lines and other equipment in the drill cellar that are essential to operating the well. Frozen material in the drill cellar is several feet thick, and at this time crews are not able to estimate how long it will take to complete the necessary thawing operations.

Two ADEC personnel stationed at the Palm Pad forward operating base are en route to the Qugruk #2 Pad for a site visit at the time of this report.


February 19, 2012

The well control team continues to thaw and evaluate equipment on the drilling rig. Crews are working to bypass the frozen drilling mud injection line to regain the capability to put mud into the well. They are also working to restore full function to the hydraulic system that operates the rig’s annular preventer. Today workers began clearing frozen water and drilling mud from the entrance to the drill floor, the drill floor itself and the drill cellar. Delineation and cleanup of the off-pad spill area remain on hold pending completion of the well control work. U.S. Fish and Wildlife Service personnel did not find evidence of a polar bear den in the spill area, but for the safety of cleanup personnel, certified bear hazers employed by Alaska Clean Seas (ACS) remain on-site round-the-clock. Two ADEC personnel remain stationed at the Palm Pad forward command post to monitor well control work and assist with site delineation and the start of cleanup. An ADEC inspector conducted a document inspection at the rig yesterday and found that all required on-site documentation was in place.


February 18, 2012

The well control team began last night and continues today to thaw equipment on the drilling rig and to evaluate its condition. The team is not yet able to estimate when the rig will be ready to return to service. Air monitoring continues to indicate safe working conditions around the rig. The start of well fluids cleanup awaits completion of the well control work. Photographs from yesterday’s fixed-wing overflight of the drilling pad show a deposit that may include drilling mud, water and cuttings from the well extending 100 feet or more from the edge of the pad. An additional area may have been misted with the same materials. Another overflight was conducted this afternoon by helicopter to provide North Slope Borough personnel and an ADEC responder a more detailed look at the spill site. Two ADEC personnel remain stationed at the Palm Pad forward command post. One will conduct a document inspection at the Qugruk #2 Pad as soon as today. Two U.S. Fish and Wildlife Service personnel are scheduled to arrive at Palm Pad at 2:00 PM today to determine whether a collared polar bear may be denning in the vicinity of the spill site.

Yesterday, in an incident unrelated to the Qugruk #2 Pad gas release, residents of Nuiqsut reported smelling natural gas in the village. The source of the odor was a vent on a building where methanethiol (also called mercaptan), a chemical used to give a characteristic odor to natural gas in the village’s gas distribution system, is stored. An Alaska Clean Seas team detected no natural gas during air monitoring conducted today in Nuiqsut at Repsol’s request. In the interest of safety, the North Slope Borough has hired a contractor to continue air monitoring in the village.


February 17, 2012

The well control team, comprised of personnel from Wild Well Control Inc., Repsol and the Alaska Oil and Gas Conservation Commission (AOGCC), continues to assess safety issues and the condition of rig equipment. AOGCC is responsible for state oversight of the well control operation. Work crews have started rigging scaffolding, tarps and heaters to thaw rig equipment, but the amount of time that will be required to restore necessary function is not known. The ambient temperature at the site today was reported to be -18°F. Three WWC personnel and an AOGCC inspector are on-site to provide technical consultation and oversight. During a flight over the spill site late yesterday, observers acquired Forward Looking Infrared (FLIR) video imagery that can be used to make a rough estimate of the spill size. ADEC personnel estimate from the imagery that the off-pad area affected by drilling mud and water is at least three acres. Another overflight departed Deadhorse at around 11:30 AM today with an ADEC responder and North Slope Borough Representative on board. Information they collected on that overflight is currently being analyzed. Four ADEC spill responders and two facility inspectors are currently based at the forward command post at Palm Pad, a staging pad approximately 26 miles east of the Qugruk #2 pad by ice road.


February 16, 2012

A team from Wild Well Control Inc. (WWC) arrived at the drilling pad at 5:30 AM today. Gas levels at the pad were low enough for the team to work safely on the pad and begin their assessment of the well. Air monitoring shows that the flow of gas from the well is minimal. WWC is working with Repsol engineers and the Alaska Oil and Gas Conservation Commission (AOGCC) on a well control plan. A planned overflight of the drilling pad to view the spill was cancelled this afternoon due to fog. An overflight will be made at the first opportunity, with an ADEC responder on board. Three other ADEC responders are en route to begin round-the-clock monitoring of the cleanup.


February 15, 2012 1930

All personnel evacuated from the pad due to safety concerns with the methane gas. A Unified Command incorporating Repsol, the U.S. Environmental Protection Agency (EPA), and ADEC, has assembled in Anchorage to manage the incident. The O’Briens Group is representing Repsol on the incident management team. Repsol has hired Wild Well Control Inc. of Houston, Texas, to assist with controlling the well, and their crew is expected to arrive on-site by 7:00 AM tomorrow. Repsol is coordinating well control efforts with the Alaska Oil and Gas Conservation Commission. Four ADEC responders are en route to the site to evaluate the incident and response. At 5:30 PM today, Repsol reported that the flow of gas from the well appeared to have nearly stopped, and gas levels near the pad were falling to safe levels.


February 15, 2012 1600

Repsol evacuated all personnel from the pad due to safety concerns associated with anticipated high levels of methane gas. An incident management team has been activated in Anchorage. Repsol has contracted the O’Briens Group to manage the incident. Repsol is researching the possibility of the Wild Well Control Company from Houston, Texas to be flown to Prudhoe Bay to assist controlling the well. Four ADEC staff will be at the incident command post established in Anchorage and one ADEC staff will arrive in Prudhoe Bay tomorrow morning to evaluate the incident and response. Alaska Clean Seas, a North Slope response company, is responding to the release.


. return to top